OpenStudy (anonymous):

what number must be added to both the numerator and the denominator of the fraction 1/5 to get a fraction equaling 3/4

OpenStudy (binary3i):

11

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

i got -11 but i dont understand why

OpenStudy (anonymous):

\[\frac{1+x}{5+x}=\frac{3}{4}\]

OpenStudy (anonymous):

then do you do 4(1+x) = 5 (3+x)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

correct nikkysterr, solve for x to get the answer :)

OpenStudy (anonymous):

yeah but i get -11

OpenStudy (anonymous):

4 + 4x = 15 + 5x

OpenStudy (anonymous):

4x = 11 + 5x

OpenStudy (anonymous):

-1x = 11

OpenStudy (anonymous):

x = -11?

OpenStudy (anonymous):

That is because it's not 4(1+x) = 5 (3+x) it should be 4(1+x)= 3(5+x)

OpenStudy (anonymous):

how about for 2/3 = 1/4

OpenStudy (anonymous):

Now you should get 11 as the answer

OpenStudy (anonymous):

\[\frac{2+x}{3+x}=\frac{1}{4}\] Same method as the last one.

OpenStudy (anonymous):

i get 4(2+x) = 1(3+x)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

but i get a -5/3

OpenStudy (anonymous):

8+4x = 3 +1x

OpenStudy (anonymous):

yes that's the answer

OpenStudy (anonymous):

4x = -5 +1x 3x = -5

OpenStudy (anonymous):

its suppose to be positive

OpenStudy (anonymous):

not always, it just says it needs to be added.